OpenStudy (anonymous):

In Radhika’s house there is a flower pot. The sum of radii of circular top and bottom of the flower pot is 140 cm and the difference of their circumferences is 88 cm. Find the diameters of the circular top and bottom

ganeshie8 (ganeshie8):

say, \(d_1\) and \(d_2\) are the diameters of circulat top and bottom

ganeshie8 (ganeshie8):

The sum of radii of circular top and bottom of the flower pot is 140 cm gives u : \(d_1 + d_2 = 2*140\)

ganeshie8 (ganeshie8):

the difference of their circumferences is 88 cm. gives u \(\pi d_1 - πd_2=88\)

ganeshie8 (ganeshie8):

two equations and two unknowns u can solve them ?

OpenStudy (anonymous):

thnx

ganeshie8 (ganeshie8):

np :) why the troll face lol

OpenStudy (anonymous):

cauz it suits my nature bro

ganeshie8 (ganeshie8):

i dont think so uhmm

OpenStudy (anonymous):

heeh ol

ganeshie8 (ganeshie8):

dont try to be something you're not :P

OpenStudy (anonymous):

you dont know me yet lol ask ikram

ganeshie8 (ganeshie8):

i trolled her too much.. she is not talking to me now :'(

OpenStudy (anonymous):

just ask about me say raj

OpenStudy (ikram002p):

-.-

ganeshie8 (ganeshie8):

w.e

OpenStudy (anonymous):

jk

ganeshie8 (ganeshie8):

ikr.. nobody can troll her... she is the biggest troll ever :|

OpenStudy (anonymous):

were jk

OpenStudy (anonymous):

ope

OpenStudy (anonymous):

dope

OpenStudy (ikram002p):

:'(

OpenStudy (anonymous):

hey gan behave yourself

ganeshie8 (ganeshie8):

im done in this thread

OpenStudy (ikram002p):

me too

ganeshie8 (ganeshie8):

??

ganeshie8 (ganeshie8):

my behavior is a direct response to urs

OpenStudy (ikram002p):

yeah right w.e

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

anything im out of here thnx gan8 i was jk

ganeshie8 (ganeshie8):

:) i can see when a non-troll *tries* to imitate a troll lol

OpenStudy (anonymous):

ah i lets see